OpenStudy (anonymous):

Which fraction shows a correct way to set up the slope formula for the line that passes through the points (4, -6) and (-2, 1)?

OpenStudy (anonymous):

(1 - (-6))/((-2) - 4) = 7/-6

OpenStudy (anonymous):

y2-y1 ----- x2-x1 -6-1 ---- 4--2 -7/6

OpenStudy (anonymous):

first find the slope, which is \[m=\frac{y_2-y_1}{x_2-x_1}\] then use almighty point - slope formula \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

both answers above are correct, slope is \[-\frac{7}{6}\]

OpenStudy (anonymous):

so now \[y-1=-\frac{7}{6}(x-(-2))\] will give you your answer